OpenStudy (anonymous):

solve 3 sin^2 x - 2 = cos x

OpenStudy (tyteen4a03):

Remember that \(sin^2(x) + cos^2(x) = 1\).

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Is 3 sin^2(x-2)=cos x

OpenStudy (tyteen4a03):

@ashwinjohn3 I believe it's 3sin^2(x) - 2.

OpenStudy (anonymous):

yea its 3 sin^2 (x) - 2 = cos (x)

hartnn (hartnn):

change sin^2 x to cos^2 x using \(\sin^2(x) + \cos^2(x) = 1\) then if you put cos^2 x = y, you'll get a quadratic in y.

OpenStudy (anonymous):

ok... then,here \[\sin ^{2} x=1-\cos ^{2} x\] \[3 (1-\cos ^{2}x )-2=3-3 \cos ^{2}x -2=1-3\cos ^{2}x\]

OpenStudy (anonymous):

thats what i get but then i cant figure out how to solve the angle :/

hartnn (hartnn):

put cos x =y, then you get a quadratic in y..

OpenStudy (anonymous):

But \[1=\cos ^{2}x+\sin ^{2} x\] =\[\sin ^{2}x+\cos ^{2}x-3\cos ^{2}x=\sin ^{2}x-2\cos ^{2}x\]

hartnn (hartnn):

?

hartnn (hartnn):

\(3 \sin^2 x - 2 = \cos x \\ 1-3\cos^2x = \cos x \\ 1-3y^2 =y. \\ 3y^2 +y-1=0 \) can you solve this quadratic ?

OpenStudy (anonymous):

yeaaa but by using calculator, the answer is in decimal point. is there any other way i can calculate the angle?

hartnn (hartnn):

you'll get same answer using any other method, that you got on calculator. whatever you got, equate it to y= ..., ... so, cos x = ... , .... then take cos^{-1} (or arccos) of those 2 values ("decimal values"), using calculator, and ou'll get 2 angles.

hartnn (hartnn):

you might be getting, -0.767 and 0.434 then 2 angles will be cos^{-1}(-0.767) and cos^{-1}(0.434) calculate those using calculator.

OpenStudy (anonymous):

hmmm okay thanks

hartnn (hartnn):

welcome :)